The coordinates of the foot of the perpendicular from the point (2,3) on the line x+y=11=0 is

Grade:12

1 Answers

Susmita
5 years ago
x+y+11=0 
Or,y=-x-11
This line has m=-1.
So the line perpendicular to it must have m=1.
equation of that line is
y=x+c
This line passes through (2,3).
3=2+c 
Or,c=1
Eq of perpendicular line is
y=x+1
Foot means point of intersection of the two lines.
Put y=x+1 in
y+x+11=0
Or,x+1+x+11=0
Or,x=-6
So y=-5
(-6.-5) is foot of the perpendicular.